OpenStudy (lgbasallote):

how do i simplify this?? \[\frac{x^2 + 3x +2}{x^2 - x - 2}\]

OpenStudy (lgbasallote):

do i factor?

OpenStudy (shubhamsrg):

yes..factoize both num and denom.

OpenStudy (shubhamsrg):

factorize*

OpenStudy (anonymous):

Yes, then perhaps they will have a common factor which you can cancel out.

OpenStudy (anonymous):

Numerator: x^2 + x + 2x + 2 = (x + 1)(x + 2) Denominator: x^2 - x - 2 = x^2 + x - 2x - 2 = (x + 1)(x - 2)

OpenStudy (lgbasallote):

so uhh some hints on how to factor this? :)

OpenStudy (anonymous):

I just did that.

OpenStudy (lgbasallote):

oh

OpenStudy (lgbasallote):

so x+1 gets canceled?

OpenStudy (anonymous):

That's riht, you'll be left with (x + 2)/(x - 2).

OpenStudy (lgbasallote):

ohhhh

OpenStudy (shubhamsrg):

in ques it should be mentioned x not equal to -1 though..

OpenStudy (anonymous):

x = -1 will give denominator as 0.

OpenStudy (lgbasallote):

uhmm could tisalso be simplfied by dividing?

OpenStudy (anonymous):

I'm not sure ... I don't think so.

OpenStudy (lgbasallote):

oh hmm